### v1

@@ 3.3 Syntactic methods
Syntactic approaches involve analysing the structure of sentences in terms of grammatical and language rules zhang2023syntactic . This contrasts with other approaches like analysing the occurrence of lexicons and interpreting textual meaning. To elaborate on the latter, while semantic methods can involve a higher level of abstraction (i.e., comprehending the hidden or implied connotations of words), syntactic measures focus on the organisation of words and phrases within a sentence without interpreting them more broadly.
Syntactic methods include part-of-speech (POS) tagging chiche2022pos , dependency parsing li2020dependency , and constituency parsing yang2020constituency , amongst others. These methods can augment efforts to accomplish several complex NLP tasks for corporate sustainability analysis, although they have only seen limited application so far. For instance, when combined with other methods like semantic analysis, text can be analysed more granularly. In other domains such as finance and economics, aspect-based sentiment analysis (ABSA) has already been achieved through POS tagging and semantic rules mao2021bridging ; consoli2022fine . Such integrated approaches can more intricately tackle task (T2) sentiment classification, by revealing sentiment insights toward aspects. To illustrate with an example, within the sentence “Company X has good emissions performance despite having poor employee wellness", ABSA allows for the aspect ‘emissions performance’ to be labelled as positive and the aspect ‘employee wellness’ to be labelled as negative. Distinguishing between the sentiment of the two aspects enables a finer analysis of sustainability text. Additionally, syntactic analysis can also be used for identifying companies and their relationships with respect to sustainability activities automatedESG_jointentity_relation . Such methodologies can be enabled by explicit syntactic attributes (i.e., POS, dependency parsing) in addition to semantic patterns. Topics can be derived from the extracted links between companies and their sustainability practices to solve task (T1) topic extraction. Additionally, these links can be further processed to evaluate corporate sustainability performance, in the spirit of task (T3) rating. Other applications of syntactic analysis include evaluating the grammatical clauses of disclosure sentences Clarkson2020 , in line with (T5) linguistic patterns, or the parsing of sustainability-related concepts, in a similar vein to the extraction of financial concepts du2023finsenticnet .
@@ 4 XNLP enhances subjectivity mitigation
Interpretability. An AI model’s capacity to provide an understanding of its mechanism sevenpillars . For example, scoring which keywords are most salient for an algorithm’s classification schuff2022human .
Explainability. An AI model’s capacity to explain why it produces an output sevenpillars . For example, explaining the reasoning steps for model decisions, as enabled by large language models jie2024interpretable .
Faithfulness. The accuracy of an AI model’s interpretability and explainability with respect to its true workings liu2022rethinking . For example, the extent to which a model’s feature salience scores and provided explanations are representative of its actual workings.
@@ 4.1 Achieving trustworthy insights
XAI capabilities allow NLP models to extend beyond classifications and predictions, providing analysts useful insights through faithful explanations and interpretations lipton2018mythos . These insights enhance NLP trustworthiness within corporate sustainability analysis, by clarifying model decisions along with the ambiguity of sustainability disclosures.
For example, figure 2 highlights how interpretability, explainability and faithfulness enhances trust in an NLP model’s classification for sustainability text. Interpretability allows an analyst to verify that appropriate and relevant text features (i.e. ‘rearing animals’ instead of ‘portfolios’), are exploited for classifying the sentence’s risk sentiment. Explainability explicates the relevance of ‘rearing animals’ to sustainability risk, further justifying the model’s classification. It explains the link between ‘rearing animals’, deforestation, methane emissions, and strained water resources, resolving the ambiguity surrounding the term ‘rearing animals’. Faithfulness ensures that the insights from interpretability and explainability are meaningful in that they accurately reflect the model’s mechanism. Put together, interpretability, explainability and faithfulness prove the robustness of the model’s decision-making process, increasing an analyst’s trust in its classification output.
Integrating these capabilities within NLP has already proven to be of significant for trust-dependent sectors like healthcare han2022hierarchical . This reinforces their broader applicability to corporate sustainability analysis, where trustworthy and credible NLP insights are also critical guix2022trustworthy . Such integration will allow NLP models to produce more actionable insights for analysts.
@@ 4.2 Inferring patterns for greenwashing research
Interpretable NLP models can afford cues on the interdependence between different features within sustainability disclosures. This can inspire researchers to theorise and subsequently validate the causal and correlation relationships between attributes relevant to corporate sustainability analysis lipton2018mythos . By doing so, our understanding of under-explored issues can be augmented, mitigating longstanding issues within the domain.
To elaborate, white-box models or intrinsically interpretable models (Naive Bayes Classifiers, Generalised Additive Models, Decision Trees etc.), allow us to grasp the decision-making process from feature inputs to classification, implying the significance of specific features and their relationships with predictions. For instance, text classification models designed with interpretability techniques can elucidate the influence of keywords for classification decisions zhang2023neurosymbolic . Developing a similar white-box model for classifying greenwashing texts allows us to study how specific text features result in a positive classification. This can offer hints on the semantic and syntactic features that characterise greenwashed texts. From a ‘machine’ based perspective, it can powerfully reveal latent data patterns easy for an analyst to miss. This fresh angle would strengthen existing greenwashing research that predominantly involves human analysis of sustainability texts de2020concepts , making greenwashing more understandable and detectable. In this fashion, interpretability can ameliorate the unreliability and ambiguity issues in the field arising from the greenwashing phenomenon.
@@ 4.3 Generalisability across frameworks and sources
Sustainability disclosures stem from different sources and are presented through diverse frameworks. As such, they can have differing linguistic and textual features, posing a challenge for NLP models to generalise. To elaborate, while firms can focus corporate sustainability disclosures on positive efforts Cho2010 , government-linked sustainability disclosures may mandate firms to disclose sustainability related risks (such as U.S. Securities and Exchange Commission 10k reports) kim2023real . A sentiment classification model trained on corporate sustainability disclosures may not adapt well to government-linked sustainability disclosures, as the latter more frequently carries negative sentiment. On top of heterogeneous sustainability data, the generalisability problem is further compounded by the lack of available datasets within the field LEE2023119726 . This reduces the diversity of data for training robust models. In light of limited data availability, interpretability methods provides an alternative means for mitigating generalisability issues.
Specifically, interpretability allows us to understand a model’s sensitivity to training data, providing insight into portions of data that cause overfitting. Recent works such as nickl2023memory detail a method to do so through a memory perturbation equation. The paper faithfully derives ‘shirt, pullover’ as classes the model is most sensitive to while training on the FMNIST dataset, and demonstrates how removing these classes can improve model generalisability performance. Moreover, it also highlights how specific samples and training epochs impact model sensitivity. While this method has been deployed for computer vision, the same principles can be adapted to NLP for sustainability analysis. Leveraging such a tool, NLP practitioners can consider removing portions of data most prone to model overfitting. This optimises training for general performance across the different sustainability disclosures. By enhancing generalisability, NLP models can effectively analyse diverse types of information associated with a firm’s sustainability efforts LEE2023119726 . This potentially compensates for incompleteness and unreliability data problems, by reducing over-reliance on specific sources that omit important details or incredible information.
@@ 4.4 Mitigating bias
By producing consistent and reproducible insights, NLP can partially address the potential bias of analysts that conduct corporate sustainability analysis. However, NLP also risks introducing its own biases if not designed transparently. To elaborate, NLP constructed with XAI capabilities can elucidate its decision-making process, allowing analysts to verify its learned features raghavan2020mitigating . While this may still incur the partiality of analyst judgement, an XAI enhanced model can conversely highlight potentially overlooked features to analysts, guiding them to produce more balanced corporate sustainability assessments. Consequently, XAI capabilities can enable a synergistic interaction between NLP models and analysts for reducing potential bias.
[…]
Acknowledgments

### v2

@@ 3.3 Syntactic methods
Syntactic approaches involve analysing the structure of sentences in terms of grammatical and language rules [89]. This contrasts with other approaches like analysing the occurrence of lexicons and interpreting textual meaning. To elaborate on the latter, while semantic methods can involve a higher level of abstraction (i.e., comprehending the hidden or implied connotations of words), syntactic measures focus on the organisation of words and phrases within a sentence without interpreting them more broadly.
Syntactic methods include part-of-speech (POS) tagging [15], dependency parsing [48], and constituency parsing [87], amongst others. These methods can augment efforts to accomplish several complex NLP tasks for corporate sustainability analysis, although they have only seen limited application so far. For instance, when combined with other methods like semantic analysis, text can be analysed more granularly. In other domains such as finance and economics, aspect-based sentiment analysis (ABSA) has already been achieved through POS tagging and semantic rules [54, 19]. Such integrated approaches can more intricately tackle task (T2) sentiment classification, by revealing sentiment insights toward aspects. To illustrate with an example, within the sentence “Company X has good emissions performance despite having poor employee wellness”, ABSA allows for the aspect ‘emissions performance’ to be labelled as positive and the aspect ‘employee wellness’ to be labelled as negative. Distinguishing between the sentiment of the two aspects enables a finer analysis of sustainability text. Additionally, syntactic analysis can also be used for identifying companies and their relationships with respect to sustainability activities [29]. Such methodologies can be enabled by explicit syntactic attributes (i.e., POS, dependency parsing) in addition to semantic patterns. Topics can be derived from the extracted links between companies and their sustainability practices to solve task (T1) topic extraction. Additionally, these links can be further processed to evaluate corporate sustainability performance, in the spirit of task (T3) rating. Other applications of syntactic analysis include evaluating the grammatical clauses of disclosure sentences [18], in line with (T5) linguistic patterns, or the parsing of sustainability-related concepts, in a similar vein to the extraction of financial concepts [28].
@@ 4 XNLP enhances subjectivity mitigation
Interpretability. An AI model’s capacity to provide an understanding of its mechanism [13]. For example, scoring which keywords are most salient for an algorithm’s classification [74].
Explainability. An AI model’s capacity to explain why it produces an output [13]. For example, explaining the reasoning steps for model decisions, as enabled by large language models [88].
Faithfulness. The accuracy of an AI model’s interpretability and explainability with respect to its true workings [50, 83, 85]. For example, the extent to which a model’s feature salience scores and provided explanations are representative of its actual workings.
@@ 4.1 Achieving trustworthy insights
XAI capabilities allow NLP models to extend beyond classifications and predictions, providing analysts useful insights through faithful explanations and interpretations [49, 56]. These insights enhance NLP trustworthiness within corporate sustainability analysis, by clarifying model decisions along with the ambiguity of sustainability disclosures.
For example, figure 2 highlights how interpretability, explainability and faithfulness enhances trust in an NLP model’s classification for sustainability text. Interpretability allows an analyst to verify that appropriate and relevant text features (i.e. ‘rearing animals’ instead of ‘portfolios’), are exploited for classifying the sentence’s risk sentiment (similar to the concepts used in self-interpretable image classification – e.g., [84]). Explainability explicates the relevance of ‘rearing animals’ to sustainability risk, further justifying the model’s classification. It explains the link between ‘rearing animals’, deforestation, methane emissions, and strained water resources, resolving the ambiguity surrounding the term ‘rearing animals’. Faithfulness ensures that the insights from interpretability and explainability are meaningful in that they accurately reflect the model’s mechanism. Put together, interpretability, explainability and faithfulness prove the robustness of the model’s decision-making process, increasing an analyst’s trust in its classification output.
Integrating these capabilities within NLP has already proven to be of significant for trust-dependent sectors like healthcare [39]. This reinforces their broader applicability to corporate sustainability analysis, where trustworthy and credible NLP insights are also critical [38]. Such integration will allow NLP models to produce more actionable insights for analysts.
@@ 4.2 Inferring patterns for greenwashing research
Interpretable NLP models can afford cues on the interdependence between different features within sustainability disclosures. This can inspire researchers to theorise and subsequently validate the causal and correlation relationships between attributes relevant to corporate sustainability analysis [49, 56]. By doing so, our understanding of under-explored issues can be augmented, mitigating longstanding issues within the domain.
To elaborate, white-box models or intrinsically interpretable models (Naive Bayes Classifiers, Generalised Additive Models, Decision Trees etc.), allow us to grasp the decision-making process from feature inputs to classification, implying the significance of specific features and their relationships with predictions. For instance, text classification models designed with interpretability techniques can elucidate the influence of keywords for classification decisions [90]. Developing a similar white-box model for classifying greenwashing texts allows us to study how specific text features result in a positive classification. This can offer hints on the semantic and syntactic features that characterise greenwashed texts. From a ‘machine’ based perspective, it can powerfully reveal latent data patterns easy for an analyst to miss. This fresh angle would strengthen existing greenwashing research that predominantly involves human analysis of sustainability texts [21], making greenwashing more understandable and detectable. In this fashion, interpretability can improve the unreliability and ambiguity issues in the field arising from the greenwashing phenomenon.
@@ 4.3 Generalisability across frameworks and sources
Sustainability disclosures stem from different sources and are presented through diverse frameworks. As such, they can have differing linguistic and textual features, posing a challenge for NLP models to generalise. To elaborate, while firms can focus corporate sustainability disclosures on positive efforts [16], government-linked sustainability disclosures may mandate firms to disclose sustainability related risks (such as U.S. Securities and Exchange Commission 10k reports) [45]. A sentiment classification model trained on corporate sustainability disclosures may not adapt well to government-linked sustainability disclosures, as the latter more frequently carries negative sentiment. On top of heterogeneous sustainability data, the generalisability problem is further compounded by the lack of available datasets within the field [47]. This reduces the diversity of data for training robust models. In light of limited data availability, interpretability methods provides an alternative means for mitigating generalisability issues.
Specifically, interpretability allows us to understand a model’s sensitivity to training data, providing insight into portions of data that cause overfitting. Recent works such as [60] detail a method to do so through a memory perturbation equation. The paper faithfully derives ‘shirt, pullover’ as classes the model is most sensitive to while training on the FMNIST dataset, and demonstrates how removing these classes can improve model generalisability performance. Moreover, it also highlights how specific samples and training epochs impact model sensitivity. While this method has been deployed for computer vision, the same principles can be adapted to NLP for sustainability analysis. Leveraging such a tool, NLP practitioners can consider removing portions of data most prone to model overfitting. This optimises training for general performance across the different sustainability disclosures. By enhancing generalisability, NLP models can effectively analyse diverse types of information associated with a firm’s sustainability efforts [47]. This potentially compensates for incompleteness and unreliability data problems, by reducing over-reliance on specific sources that omit important details or incredible information.
@@ 4.4 Mitigating bias
By producing consistent and reproducible insights, NLP can partially address the potential bias of analysts that conduct corporate sustainability analysis. However, NLP also risks introducing its own biases if not designed transparently. To elaborate, NLP constructed with XAI capabilities can elucidate its decision-making process, allowing analysts to verify its learned features [69]. While this may still incur the partiality of analyst judgement, an XAI enhanced model can conversely highlight potentially overlooked features to analysts, guiding them to produce more balanced corporate sustainability assessments. Consequently, XAI capabilities can enable a synergistic interaction between NLP models and analysts for reducing potential bias.